Please help me solve this word problem: You invest $100 and expect your money to grow 8% each year. About how many years will it take for your investment to double?
What function models the growth of your investment?
How can you use a table to find the approximate amount of time it takes for your investment to double?
How can you use a graph to find the approximate amount of time it takes for your investment to double?

Answer by Boreal(15235)

The rule of 72 would say 9 years. (doubling time in years=72/interest rate in per cent.
Exactly, it is P=Po(1+.08)^t
Po is start, P is end
P/Po therefore equals 2.
2=1.08^t
ln2=t ln (1.08)
t=9.01 years.
table
1------1.08
2------1.08^2=1.1664 etc. until one reaches 2.
This can be used graphically, and one could draw a curve through the points until 2 is reached for the amount of money.